Systems Staff Meeting Notes 6/1/98

Systems Staff Meeting Notes
June 1, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Sharon Henry, Bobbie Parker, Suzy Shaw, Leota Whitfield

Announcements
  1. Destry turned in his resignation letter, effective June 18. No temps will be hired for his position. We will advertise & try to get a full-scale hire.
  2. Year-end buying is in full force - Barbara would like to close out by the end of this week.
  3. The state is allowing personal orders for PCs to be ordered through the state - may need to watch out for individual orders coming in with business ones.
  4. Water in the restrooms is brown in color - has been reported to John Humphrey.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Count on Cows on order up to 45 - some are special orders.
  2. Two Cows have appeared behind the curtain.
  3. Four Okidata printers have been received.
Networking
  1. MSL is scheduled for the switch install next week.
  2. AFA & Music can be added easily. The IPs & Gateways will change. Rich will check on how the Circulation access is defined.
  3. After the switch in West, a problem occurred with people being kicked off of SNA sessions. We think it was a local problem - the uplink port was misconfigured. Bill changed it on Friday afternoon, no problems have been reported since. The uplink was set to half-duplex, it has been changed to full-duplex. SNA is not routed protocol - timing problems. A sniffer is currently activated at NERDC. If anyone is aware of problems with SNA sessions, it is crucial they are reported to Bill.
  4. Another recurring problem after the switch install has been the phantom disappearance of the Smathers server. NERDC activated a sniffer on Friday in AFA - could see packet & header was correct - could see lead & end, but not Smathers. The problem occurred in MSL this morning. Looks like it may be CORE network problem. Filters were added to Smathers, cut advertisers down to less than half.
  5. Will need to remind people of their trouble shooting skills - they need to check out problems before calling them in.
  6. Smathers hiccuped Wednesday AM. Printers were fading in & out of Novell. We have removed Novell print server queues - all should be on NT machines. No one should be pointing to Smathers for print. Unless someone is utilizing CDs, Novell access should be stripped.
  7. CD Tower should be fine. Should be accessible to all. Rich will check on their accessibility from MSL, Friday Census was not available. Some may need updates.
  8. SAP filters will be added to Music & AFA.
  9. MSL has thinnet - the switch has an AUI port which can handle it. We will get Gresh to yank the thinnet & replace it. Two machines are PS2s used for Dialog - wont handle UTP. Need to find some other way to get the service - whoever is closest to Dialog needs to find improvement.
  10. Well be working on the diagramming of the plugs & connections in MSL this week.
  11. There is talk of using leftover FCLA money for purchasing more switching hubs.
  12. Preservation - there is still a problem in UF Purchasing over sole source - money should at least be encumbered by the end of this fiscal year.
  13. MSL - wall is up, still no furniture.
Software Projects Status
  1. Sharon reports that Evelyn has been getting OBDC errors. Problem with the SQL server on NT4 - Winston fixed it by adding a patch to the server. Now it tries to write duplicates. She is unable to input records. Sharon is unable to sign into NT4 SQL.
  2. Personnel status - Sharon is working on the permanent side.
  3. Mellon - July 1 Systems will take on interfacing the database to images.
  4. Facilities - delivered Quickbook Pro to them - Library Office will be using it for Foundation work.
  5. Timer Software - Brian has spoken with John. He wants an email timer - this is a change from what was requested - now its two distinct projects. Need to continue negotiations on what is possible vs. what is desired.
  6. Bill to meet with PSITE to further discuss log ins.
DP Coordinator
  1. Debra received a notice from Compustat (via Peter McKay) requesting the return of tapes checked out on a year-only lease. Originals have been returned, copies purged.
Calendar
  1. Bobbie out June 8 - 19.
  2. Debra out half day June 16, all day June 17.
  3. Winston out June 15-24.
  4. ALA in Washington, D.C. is June 24 - 30.
  5. Rich will be gone to ALA/vacation June 24-July 15.
  6. Independence Day observed July 3.
  7. Brian and Lee will be out July 6.
  8. Debra out July 7 - 10.
  9. Bobbie out July 13 - 17
  10. Brian out July 16 & 17.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu